28. Factor completely. 21x2 + 33x + 12

Answer by checkley71(8403)

(3X+3)(7X+4)=0 OR 3X+3=0 OR 3X=-3 OR X=-3/3 OR X=-1 & 7X+4=0 OR 7X=-4 OR X=-4/7